Max and Sam have the same size lawn to mow max can finish in 24 min while Sam finished in 36 min
At what time did Sam have twice as much to mow as Max

Answer by richwmiller(17219)

s=1-t/24,
m=1-t/36,
m=2s

2*(1-x/24) =(1-x/36)
2-x/12 = 1-x/36
2-x/12 = 1-x/36
1-x/12=-x/36
1=-x/36+x/12
1=-x/36+3x/36
1=2x/36
1=x/18
x=18